OpenStudy (anonymous):

find dy/dx by implicit differentiation sqrt (x+y) = (5+x^2y^2)

OpenStudy (anonymous):

will give medal & fan ... but i want to understand the steps too...

OpenStudy (anonymous):

@iambatman @phi @PaxPolaris @mathsails

OpenStudy (anonymous):

@Yttrium

OpenStudy (anonymous):

someone help me :'(

OpenStudy (kainui):

\[\LARGE \sqrt {x+y} = (5+x^2y^2)\] Ok try your best to take the derivative and I'll help you.

OpenStudy (kainui):

Well I'm not going to wait around all day, I can help you out and I know the answer but I won't help someone who won't help themselves first.

OpenStudy (zephyr141):

i have something.... \[\frac{ 1 }{ 2\sqrt{x+y} }*(1+\frac{ dy }{ dx })=2x*2y*\frac{ dy }{ dx }\]

OpenStudy (anonymous):

except that was wrong

OpenStudy (phi):

there is a bit of algebra for this one. How far did you get? did you find the derivative of the left side? what did you get ?

OpenStudy (anonymous):

1/2 (x+y)^-1/2

OpenStudy (anonymous):

and at the end of that I would add d/dx

OpenStudy (anonymous):

that's what i'm trying to get to one side to solve for

OpenStudy (anonymous):

I have to go on the skytrain now

OpenStudy (anonymous):

:'(

OpenStudy (phi):

\[ \frac{d}{dx} u^\frac{1}{2} = \frac{1}{2}u^{-\frac{1}{2}} \frac{d}{dx} u \] you have the first part, but you need the second part \[ \frac{d}{dx} u = \frac{d}{dx} (x+y) \]

OpenStudy (anonymous):

this is due in 40 mins oh dear lord

OpenStudy (phi):

the left side is \[ \frac{1}{2 \sqrt{x+y}} \cdot (1 + \dot{y} )\]

OpenStudy (phi):

the right side is x^2 d/dx (y^2) + y^2 d/dx(x^2)

OpenStudy (phi):

solve for \( \dot{y} \)

OpenStudy (phi):

for the right side you should get \[ 2 x^2 y \ \dot{y} + 2 x y^2 \]

OpenStudy (phi):

now it's algebra, solve for \( \dot{y} \) \[ \frac{1}{2 \sqrt{x+y}} \cdot (1 + \dot{y} ) = 2 x^2 y \ \dot{y} + 2 x y^2 \]